OpenStudy (anonymous):

If sin(x)-cos(x)=1/2,find sin(x)+cos(x)

OpenStudy (ahsome):

@hartnn we need your help!

OpenStudy (ahsome):

Is this the equation:\[sin\theta-cos\theta=\frac{1}{2}\]

OpenStudy (anonymous):

yes

OpenStudy (ahsome):

Do they give you \(\theta\)?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

any idea bro?

OpenStudy (ahsome):

I am trying, just can't figure it out...

OpenStudy (anonymous):

same here

OpenStudy (ahsome):

What could we do to change \(sin\theta-cos\theta\) to \(sin\theta+cos\theta\)?

OpenStudy (anonymous):

well,one could use the \[\left( a-b \right)^{2 } \] identity

OpenStudy (ahsome):

How?

hartnn (hartnn):

Blai is correct with \((\sin x - \cos x )^2 \) you will get the value if 2 sin x cos x that you can use in \((\sin x + \cos x )^2 \) and just take the square root :)

hartnn (hartnn):

remember that \(\sin^2 x +\cos^2 x =1\)

OpenStudy (ikram002p):

nice !

OpenStudy (anonymous):

(sinx-cosx)^2=(1/2)^2

OpenStudy (ahsome):

Great Job!

hartnn (hartnn):

yes chad, let blai try out further steps :)

OpenStudy (anonymous):

sin^2x-2sinxcosx+cos^2x=1/4

OpenStudy (anonymous):

sorry

hartnn (hartnn):

no problem :)

OpenStudy (anonymous):

go ahead blai the platform is yours

OpenStudy (ahsome):

The Questions the Stage, the mathematicians merely investigators.

OpenStudy (ahsome):

@BlaiRe, are you still here ;)

OpenStudy (tkhunny):

This is all fine if \(\sin(x) \ge \cos(x)\). What if it isn't? It might be interesting to exploit these: \(\sin(x) - \cos(x) = \sqrt{2}\sin(x-\pi/4)\) \(\sin(x) + \cos(x) = \sqrt{2}\sin(x+\pi/4)\)

hartnn (hartnn):

if sin x < cos x, wouldn't sin x - cos x be negative ? sin x-cos x = 1/2 implies sin x is greater than cos x by 1/2

OpenStudy (tkhunny):

True, but when squaring, you have introduced other solutions. Are you SURE you have eliminated them along the way?

OpenStudy (anonymous):

I'm here now

OpenStudy (anonymous):

OK,here's what I've got after careful analysis of the above replies \[\left( \sin \theta -\cos \theta \right)^{2} =\frac{ 1 }{ 4 } \] \[1-2\sin \theta \cos \theta =\frac{ 1 }{ 4 }\] \[-2\sin \theta \cos \theta =\frac{ 1 }{ 4 } -1 \] \[\frac{ 1-4 }{ 4 }\] \[\frac{ -3 }{ 4 } \] \[2\sin \theta \cos \theta =\frac{ 3 }{ 4} \] \[\left( \sin \theta +\cos \theta \right)^{2} = \sin ^{2} +\cos ^{2} +2\sin \theta \cos \theta \] \[1+\frac{ 3 }{ 4 }\] \[\frac{ 7 }{ 4 }\] \[\sin \theta +\cos \theta =\frac{ \sqrt{7} }{ 2 }\]

hartnn (hartnn):

thats correct :) you would have got the same answer if sin x-cos x was -1/2

OpenStudy (tkhunny):

That's not quite correct. There are infinitely many solutions.